OpenStudy (anonymous):

if sinx = 1/3 what does sin^2(x) =

OpenStudy (anonymous):

\[\sin x=\frac{1}{2}~~\Rightarrow~~\left(\sin x\right)^2=\left(\frac{1}{3}\right)^2=\cdots\]

OpenStudy (anonymous):

A note on notation: \(\sin^2x=(\sin x)^2\).

OpenStudy (anonymous):

ok thanks thats what i thought and Root(sinx^\[\sqrt{\sin ^{2}} = \left| sinx \right| correct?\]

OpenStudy (anonymous):

Yes